OpenStudy (kofi):

A ball is thrown directly downward with an initial speed of 8.00 m/s, from a height of 30.0 m. After what time interval does it strike the ground?

OpenStudy (mathmate):

Advice Please post in the appropriate section, or else helpers for the section will have to look all over for questions. If no answer in the appropriate section, you can solicit help elsewhere by linking to the original question. hint: use the kinematics equation S=ut+(1/2)at^2 S=-30 m distance travelled (negative downwards) u=-8 m/s initial velocity, (negative downwards) t=time in seconds (to be solved for) a=-9.81 m/s^2 acceleration due to gravity, negative downwards) Solve for t as the time required, reject negative root (for time).

OpenStudy (kofi):

alright thank you

OpenStudy (mathmate):

you're welcome! :)

OpenStudy (ravneet):

Use V^2=U^+2AS to find final velocity and than apply V=U+AT where V=final velocity U=Initial Velocity(-8) A=Acc. due to gravity(-10 approx) S=displacement(-30) T=time of motion [ all are negative as I have taken downward direction as negative and Upward direction as Positive . you can change their sign according to your convience ] NOW FIND time SO NO NEED TO SOLVE QUADRATIC EQUATION